TestReport Reference Manual

1.0

Generated by Doxygen 1.3-rc3

Mon Jun 14 17:22:24 2004

CONTENTS  1

Contents 1 Run Test Report

Author: {automatically generated} 1.0.1 Summary There are 5273 events in /nfs/farm/g/glast/u01/svac workshop/em - v1r0401p9/mc/TwoTowers/gamma/digi/gamma 100 3.5 0 -300 50 900 - 10000 0 digi.root. 1.0.2 Trigger

Figure 1: Trigger distribution

1 Run Test Report  2

1.0.3 Strip hit info

Figure 2: Number of hit strips in tower 08

Generated on Mon Jun 14 17:22:24 2004 for TestReport by Doxygen

1 Run Test Report  3

Figure 3: Number of hit strips in tower 09

Generated on Mon Jun 14 17:22:24 2004 for TestReport by Doxygen

1 Run Test Report  4

Figure 4: Number of hit layers in tower 08

Generated on Mon Jun 14 17:22:24 2004 for TestReport by Doxygen

Table 1: Number of events with different numer of TKR digis. There are 5268 events with 3 in a row TKR trigger.

Number of TKR digis Number of TKR events percentage
0  0  0%
1  0  0%
2  0  0%
3  0  0%
4  0  0%
5  0  0%
≥ 6  5268  100%

Figure 5: Number of hit layers in tower 09

Bad Strips:

• There are 0 events with strip ID outside the range from 0 to 1535.

• There are 0 events with more than 63 strips per GTRC.

1.0.4 Time Over Threshold TOT info:

• There are 189 events with saturated TOT values.

• There are 1299 events with zero TOT on one plane but nonzero number of strip hits in that plane.

• There are 0 events with nonzero TOT on one plane but no strip hits in that plane.

Figure 6: TOT distribution